Please help me solve this:
Box 1 contains 2 red balls and 1 blue ball. Box 2 contains 3 blue balls and 1 red ball. A coin is tossed and if it falls heads up box 1 is selected and a ball is drawn. If the coin lands tails up box 2 is selected and a ball is drawn. Find the probability of selecting a red ball?

Answer by sudhanshu_kmr(1152)


probability of selecting a red ball = 1/2 * 2/3 + 1/2 * 1/4

= 1/3 + 1/8
= 11/24